OpenStudy (anonymous):

x-y+1=0 find slope pleaseee

OpenStudy (anonymous):

Solve for y, x-y+1=0 (x-y+1)-1=-1 (x-y)-x=-1-x -y/-1=(-1-x)/-1 y=1+x

OpenStudy (anonymous):

is that the slope ?

OpenStudy (anonymous):

y=mx+b where m is the slope (slope intercept form) your question is in general form x-y+1=0 -y=-x-1 divide by -1 to change y sign y=x+1 your slope is 1/1 or 1

OpenStudy (anonymous):

THANKS MIKE